OpenStudy (anonymous):

Find the shortest distance from the line y=3x+5 to the origin

OpenStudy (anonymous):



OpenStudy (anonymous):

3x-y+5=0 Perpendicular Distance from Orgin = |c| / sqrt (a^2+b^2)

OpenStudy (anonymous):

anyone here ..

OpenStudy (anonymous):

Ax+By + C = 0 Hope nw u can solve..._)

OpenStudy (anonymous):

dont really get it

OpenStudy (anonymous):

\[Distance = \frac{ |c| }{ \sqrt{a^2+b^2} }\]

OpenStudy (anonymous):

@Yahoo! what formula is this..? it seems like i never learn it before

OpenStudy (anonymous):

Lol...This is a Formula..)

OpenStudy (anonymous):

This may help u... http://answers.yahoo.com/question/index?qid=20081217164102AAcolPs

OpenStudy (anonymous):

really ok thanks..

OpenStudy (anonymous):

Welcome...